OpenStudy (anonymous):

Solve 3|2x + 2| − 7 = 5 for x

OpenStudy (anonymous):

3|2x + 2| − 7 = 5 3|2x + 2| = 5 +7 2x+2=12/3 2x=4-2 x=1